I need help plz.. Solve by the substitution method. What is the solution?
3x - 2y =-6
5x + 24 =y

Answer by stanbon(57352)

I need help plz.. Solve by the substitution method. What is the solution?
3x - 2y =-6
5x + 24 =y
------------------
Subtitute for "y" in the 1st equation and solve for "x".
---
3x - 2(5x+24) = -6
3x - 10x - 48 = -6
-7x = 42
x = -6
-----
Substitute into y = 5x+24 to solve for "y":
y = 5*-6 + 24 = -6
--------------------------
Cheers,
Stan H.